### v1

# Ramsey Numbers in Kneser Graphs

Abstract
We define the $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r}(s,t)$ as the minimum integer $n$ such that every red/blue edge-coloring of the Kneser graph $\textrm{KG}(n,r)$ contains a red $s$-clique or a blue $t$-clique. We obtain general bounds on the numbers $R^{\textrm{KG}}_{r}(s,t)$, and make progress on two related Ramsey-type problems, one raised by Holmsen, Hrusak, and Roldán-Pensado, and the other posted by Pálvölgyi.
1 Introduction
The Ramsey number $R(s,t)$ is the minimum integer $n$ such that any red/blue edge-coloring of the complete graph $K_{n}$ contains either a red $K_{s}$ or a blue $K_{t}$. Ramsey numbers and their many different variations have been extensively studied, and they form one of the main pillars of extremal combinatorics. In this paper we consider a variation on Ramsey numbers in which the host graph is a Kneser graph rather than the complete graph. The Kneser graph $\textrm{KG}(n,r)$ is the graph whose vertex set is the set $n]\choose r$} of all $r$-subsets of $[n]$, and two $r$-sets form an edge if they are disjoint. We define the Kneser Ramsey numbers as follows.
Definition 1.
Let $r\geq 1$ be an integer. The $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r}(s,t)$ is the minimum integer $n$ such that every red/blue edge-coloring of $\textrm{KG}(n,r)$ contains a red $K_{s}$ or a blue $K_{t}$.
Note that $R^{\textrm{KG}}_{1}(s,t)=R(s,t)$, so this definition generalizes the notion of Ramsey numbers.
Our motivation to investigate Kneser Ramsey numbers originated with the following fascinating Ramsey-type question, which was asked by Michael Hrusak for $\alpha=1/2$ and then by Andreas Holmsen and Edgardo Roldán-Pensado for all $\alpha$ [4]:
Problem 2 (Holmsen - Hrusak - Roldán-Pensado).
Fix $\alpha\in(0,1)$ and an integer $t>1$. Does there exist a large enough $n$ (depending on $\alpha$ and $t$) such that in every red/blue edge-coloring of the complete graph on vertex set $V={[n]\choose\alpha n}$ there is a monochromatic clique on vertex set $U\subset V$, with the property that the vertex-covering number of the hypergraph $([n],U)$ is at least $t$?
Although Problem 2 is not about Kneser graphs, one observes that a positive answer will follow if in every red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ there is a monochromatic clique of size $t$. That is, if $R^{\textrm{KG}}_{\alpha n}(t,t)\leq n$ for some $n$, then we get a positive answer to Problem 2.
However, Problem 2 is interesting even when $\textrm{KG}(n,\alpha n)$ is empty. In fact, Holmsen and Roldán-Pensado were most interested in the special case when $\alpha=k/(2k-1)$ and $t=2$, as formulated below.
Problem 3.
Let $V={[2k-1]\choose k}$. Is it true that for some large enough $k$, in every red/blue edge-coloring of the complete graph on vertex set $V$ there is a monochromatic triangle $ABC$ with $A,B,C\in V$ and $A\cap B\cap C=\emptyset$?
Perhaps a more approachable problem is the following.
Problem 4.
Determine the smallest number $\beta>1$ such that for some large enough $k$, in every red/blue edge-coloring of the complete graph on vertex set $V={[\beta k]\choose k}$ there is a monochromatic triangle $ABC$ with $A,B,C\in V$ and $A\cap B\cap C=\emptyset$.
It is easy to see that $\beta\leq 3$. Indeed, let $S_{1},\dots,S_{6}\in{[3k]\choose k}$ be six sets with the property that for every $\in[3k], tt$ belongs to exactly two of the sets $S_{1},\dots,S_{6}$. (It is easy to construct such sets $S_{1},\dots,S_{6}$: take the six distinct sets in two disjoint perfect matchings in the $k$-uniform hypergraph on vertex set $[3k]$.) Now consider a red/blue edge-coloring of the complete graph on vertices $S_{1},\dots,S_{6}$. Since $R(3,3)=6$, this graph contains a monochromatic triangle $S_{i}S_{j}S_{k}$, and $S_{i}\cap S_{j}\cap S_{k}=\emptyset$ for every distinct $i,j,k\in[6]$.
In this paper we make progress on Problems 2 and 4 in two different ways. First, in Section 2 we improve the upper bound on $\beta$ to $7/3$.
Theorem 5.
Let $k\geq 12$ be an integer divisible by 6. Then in every red/blue edge-coloring of the complete graph on vertex set $V={[\frac{7}{3}k]\choose k}$ there is a monochromatic triangle $ABC$ with $A,B,C\in V$ and $A\cap B\cap C=\emptyset$.
Second, in the same section, we give a positive answer to Problem 2 when $\alpha<1/4$ and $t=3$.
Theorem 6.
For any $0<\alpha<\frac{1}{4}$, there exists some large enough $n$ such that any red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
Note that for $\alpha=\frac{1}{4}$, for $n=8$ there exists a red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ with no monochromatic triangle, as implied by Theorem 12.
We then continue with a systematic study of Kneser Ramsey numbers for various parameters $r,s,t$. The following trivial upper bound on the $r$-Kneser Ramsey number follows from Ramsey’s theorem:
$R^{\textrm{KG}}_{r}(s,t)\leq R(s,t)\cdot r.$
Indeed, when $n\geq R(s,t)\cdot r$, the graph $\textrm{KG}(n,r)$ contains a clique of size $R(s,t)$, and any red/blue edge-coloring of it contains a red $K_{s}$ or a blue $K_{t}$. This in particular shows the existence of the number $R^{\textrm{KG}}_{r}(s,t)$. Note also that if $n<kr$, then $\textrm{KG}(n,r)$ does not contain a clique of size $k$, and thus we have a trivial lower bound
$R^{\textrm{KG}}_{r}(s,t)\geq r\cdot\max\{s,t\}.$
In order to prove Theorem 6, we prove an auxiliary result (Theorem 17) that in particular implies the following bound on $R^{\textrm{KG}}_{r}(s,t)$.
Theorem 7.
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). We have
$R^{\textrm{KG}}_{r}(s,t)\leq(M+1)r+m-1.$
In particular,
$R^{\textrm{KG}}_{r}(s,s)\leq(R(s,s-1)+1)r+R(s,s-1)-1.$
In Section 3 we establish improved upper bounds on $R^{\textrm{KG}}_{r}(s,t)$ for certain families of $s,t$.
Theorem 8.
For $s\geq 4$, we have
$R^{\textrm{KG}}_{r}(s-1,s)\leq rR(s-2,s)+2(R(s-2,s-1)-1)+2r.$
Theorem 9.
For $t\geq 5$, we have $R^{\textrm{KG}}_{2}(3,t)\leq 2(R(3,t-2))+2t+1.$
We then derive a general lower bound, and an improved lower bound for $s=t=3$ and $r\geq 3$.
Proposition 10.
For $r\geq 1$ and $s,t\geq 2$, we have $R^{\textrm{KG}}_{r}(s,t)\geq R(s,t)+2r-2.$
Theorem 11.
For $r\geq 2$, we have $R^{\textrm{KG}}_{r}(3,3)\geq 3r+2.$
Using a computer code, we also obtain an exact value on $R^{\textrm{KG}}_{2}(3,3)$. The computational proof of the next theorem is discussed in Subsection 3.3.
Theorem 12.
We have $R^{\textrm{KG}}_{2}(3,3)=9$ and $R^{\textrm{KG}}_{3}(3,3)\leq 13$.
Some of our results are summarized in Table 1 (for $r=2$) and Table 2 (for $r=3$). In each cell of the tables, the first number listed is the lower bound and the second is the upper bound. The subscripts on the numbers are the theorems in which we obtain the bound. For example, we have $11\leq R^{\textrm{KG}}_{2}(3,4)\leq 16$, where the lower bound follows from Theorem 10 and the upper bound is proven in Theorem 8. If there is only one number, then it is an exact value. Note that $R^{\textrm{KG}}_{r}(s,t)=R^{\textrm{KG}}_{r}(t,s)$.
Table 1: Bounds on $R^{\textrm{KG}}_{2}(s,t)$.
$s$ $t$ $t=3$ $t=4$ $t=5$ $t=6$
$s=3$ 912 1110, 168 1610, 239 2010, 319
$s=4$ 2010, 287 2710, 488 3810, 697
$s=5$ 4510, 767 6010, 1348
$s=6$ 10410, 2627
Table 2: Bounds on $R^{\textrm{KG}}_{3}(s,t)$.
$s$ $t$ $t=3$ $t=4$ $t=5$ $t=6$
$s=3$ 1111, 1312 1310, 247 1810, 347 2210, 507
$s=4$ 2210, 387 2910, 707 4010, 957
$s=5$ 4710, 1027 6210, 1877
$s=6$ 10610, 3507
We conclude this paper with some progress on a Ramsey-type question that was posted by Dömötör Pálvölgyi on mathoverflow in 2016 [6]. Here an induced $\textrm{KG}(3r,r)$ is a subgraph of $\textrm{KG}(n,r)$ induced on vertex set $S\choose r$}, where $S\subset[n]$ is a set of size $3r$.
Problem 13 (Pálvölgyi [6]).
For $r>1$, determine the smallest $N\geq 3r$ such that in any red/blue edge-coloring of $\textrm{KG}(N,r)$ there exists either a red induced $\textrm{KG}(3r,r)$ or a blue triangle.
In Section 4 we prove a lower bound, depending on Ramsey numbers, for the number $N$ in Pálvölgyi’s problem. For integers $r,s,t$, let $R_{r}(s,t)$ be the minimum integer $n$ such that in any red/blue coloring of the edges of the complete $r$-uniform hypergraph on $n$ vertices $K^{r}_{n}$, there is a red $s$-clique or a blue $t$-clique.
Theorem 14.
We have $N\geq R_{r}(3r,3r)$. That is, for $n=R_{r}(3r,3r)-1$, there is a red/blue coloring of $\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
For example, for $r=2$ we get $N\geq 102$, as $R_{2}(6,6)=R(6,6)\geq 102$ [3].
Many of our proofs use a theorem of Alon, Frankl, and Lovász [1].
Theorem 15 (Alon - Frankl - Lovász [1]).
Let $k_{1}\geq k_{2}\geq\dots\geq k_{t}\geq 2$ and suppose $n\geq k_{1}r+\sum_{i=2}^{t}(k_{i}-1)$. Let $F_{1},F_{2},\dots,F_{t}$ be a partition of $\binom{[n]}{r}$. Then, for some $i\in[t]$, $F_{i}$ contains $k_{i}$ pairwise disjoint sets.
Observe that when $k_{1}=k_{2}=\cdots=k_{t}=2$, Theorem 15 recovers the chromatic number of the Kneser graph. When $t=2$, we apply a simpler form.
Theorem 16 (Alon - Frankl - Lovász [1]).
Let $n\geq(r+1)k-1$ and let $F_{1},F_{2}$ be a partition of $\binom{[n]}{r}$. Then, either $F_{1}$ or $F_{2}$ contains $k$ pairwise disjoint sets.
2 The Holmsen - Hrusak - Roldán-Pensado problem
We start with the proof of Theorem 5, showing that $7/3$ is an upper bound on $\beta$ in Problem 4.
Proof of Theorem 5. Let $V=[3k/2]\cup T$, where $T$ is a set of $5k/6$ vertices disjoint from $[3k/2]$. Then $|V|=7k/3$. Suppose we have a red/blue edge-coloring of the complete graph $G$ on $V\choose k$}.
Consider the triangle $S_{1}S_{2}S_{3}$ in $G$, where $S_{1}=\{1,\dots,k\}$, $S_{2}=\{1,\dots,k/2\}\cup\{k+1,\dots,3k/2\}$, and $S_{3}=\{k/2+1,\dots,3k/2\}$. Since $S_{1}\cap S_{2}\cap S_{3}=\emptyset$, if $S_{1}S_{2}S_{3}$ is monochromatic we are done. Otherwise, it must have a red edge and a blue edge. Without loss of generality, $S_{1}S_{2}$ is red, and $S_{1}S_{3}$ is blue. Note that $S_{1}\cap S_{i}\subset[k]$ for all $i$.
Let $T_{1},T_{2},T_{3}$ be pairwise disjoint subsets of $\{k+1,\dots,3k/2\}$ of size $k/6$ each. Let $F_{i}=T\cup T_{i}$ for $i\in[3]$.
Assume first that the triangle $F_{1}F_{2}F_{3}$ is monochromatic. If it is monochromatic blue, then consider the induced graph of $G$ on vertex set $\{S_{1},S_{2},F_{1},F_{2},F_{3}\}$. Observe that if either $S_{1}$ or $S_{2}$ has at least two blue edges to vertices in $\{F_{1},F_{2},F_{3}\}$, then we get a blue triangle of the form $S_{i}F_{j}F_{l}$, and since $S_{i}\cap F_{j}\cap F_{l}=S_{i}\cap T=\emptyset$, we are done. Otherwise, both $S_{1}$ and $S_{2}$ have at least two red edges to vertices in $\{F_{1},F_{2},F_{3}\}$, so by the pigeonhole principle we get a red triangle of the form $S_{1}S_{2}F_{i}$. But now $(S_{1}\cap S_{2}\cap F_{i})\subset([k]\cap F_{i})=\emptyset$, and we are done. If the triangle $F_{1}F_{2}F_{3}$ is monochromatic red, the argument is symmetric when we consider instead the induced graph on vertex set $\{S_{1},S_{3},F_{1},F_{2},F_{3}\}$.
So suppose the triangle $F_{1}F_{2}F_{3}$ is not monochromatic. Then without loss of generality $F_{1}F_{2}$ is a red edge and $F_{1}F_{3}$ is a blue edge.
Consider the induced graph on vertex set $\{H_{1},\dots,H_{6}\}$, where $H_{i}=[k-1]\cup\{k+i-1\}$ for $i\in[6]$. Since $R(3,3)=6$, this graph has a monochromatic triangle, say $H_{1}H_{2}H_{3}$. Note that $H_{i}\cap H_{j}=[k-1]$ for $i\neq j$. Moreover, since $k/2\geq 6$ we have $H_{i}\cap T=\emptyset$ for all $i\in[6]$.
If the triangle $H_{1}H_{2}H_{3}$ is monochromatic blue, then consider the induced graph of $G$ on vertex set $\{F_{1},F_{2},H_{1},H_{2},H_{3}\}$. Observe that if $F_{1}$ or $F_{2}$ has at least two blue edges to vertices in $\{H_{1},H_{2},H_{3}\}$, then we get a blue triangle of the form $F_{i}H_{j}H_{l}$, and since $F_{i}\cap H_{j}\cap H_{l}=F_{i}\cap[k-1]=\emptyset$, we are done. Otherwise, both $F_{1}$ and $F_{2}$ have at least two red edges to vertices in $\{H_{1},H_{2},H_{3}\}$, so by the pigeonhole principle we get a red triangle of the form $F_{1}F_{2}H_{i}$. But now $(F_{1}\cap F_{2}\cap H_{i})\subset(T\cap H_{i})=\emptyset$, and again we are done.
Finally, if the triangle $H_{1}H_{2}H_{3}$ is monochromatic red, we have a symmetric argument by considering instead the induced graph on vertex set $\{F_{1},F_{3},H_{1},H_{2},H_{3}\}$. ∎
Our next goal is to prove Theorem 6, which gives a positive answer to Problem 2 in the case $\alpha<1/4$ and $t=3$. In order to prove the theorem, we need the following auxiliary result.
Theorem 17.
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). Let $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$. Then, there exists some $n\in\mathbb{N}$ such that $\alpha n\in\mathbb{N}$ and $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
Proof of Theorem 17.
Fix $r\geq 2$ and set Nr=Mr+m1N-r=Mr+m-1. First, we will show that in any red/blue coloring of $G=\textrm{KG}(N,r)$, there is a red $K_{s}$ or a blue $K_{t}$. Suppose for contradiction that there is some coloring $c:E(G)\to\{R,B\}$ of $G$ without a red $K_{s}$ or blue $K_{t}$. Let $v=[N]\setminus[N-r]\in V(G)$. Then, $\binom{[N-r]}{r}=\binom{[Mr+m-1]}{r}$ can be partitioned into $N_{R}(v)$ and $N_{B}(v)$. Since $c$ contains no red $K_{s}$ and no blue $K_{t}$, then $G[N_{R}(v)]$ contains no red $K_{s-1}$ and no blue $K_{t}$. Similarly, $G[N_{B}(v)]$ contains no red $K_{s}$ and no blue $K_{t-1}$. In particular, $G[N_{R}(v)]$ contains no $K_{M}$ and $G[N_{B}(v)]$ contains no $K_{m}$. However, by Theorem 16, setting $F_{1}=N_{R}(v)$ and $F_{2}=N_{B}(v)$, either $N_{R}(v)$ contains a clique of size $M$ or $N_{B}(v)$ contains a clique of size $m$, a contradiction.
For the general case, fix $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$. Choose $r\in\mathbb{Z}$ such that $\frac{r}{(M+1)r+m-1}\geq\alpha$. Set Nr=Mr+m1N-r=Mr+m-1 and $n=\frac{r}{\alpha}$. Then, by choice of $r$, observe that $N\leq n$. Therefore, since $\textrm{KG}(N,\alpha n)\subseteq\textrm{KG}(n,\alpha n)$, any red-blue coloring of $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$. ∎
Note that the general bound on $R^{\textrm{KG}}_{r}(s,t)$ in Theorem 7 is proven in the first paragraph of the proof of Theorem 17 above.
We are now ready to prove Theorem 6.
Proof of Theorem 6.
We apply Theorem 17 with $s=t=3$. This gives $M=m=R(2,3)=3$. For any $0<\varepsilon<\frac{1}{4}$, we have $\alpha=\frac{1}{4}-\varepsilon$, and by Theorem 17 there exists some $n\in\mathbb{N}$ with $\alpha n\in\mathbb{N}$ such that any red/blue coloring of the edges of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle. ∎
3 General bounds on Kneser Ramsey numbers
In this section we prove Theorems 8, 9, 11, and 12, and Proposition 10.
3.1 Upper bounds
Proof of Theorem 8.
We would like to show that for $s\geq 4$,
$R^{\textrm{KG}}_{r}(s-1,s)\leq rR(s-2,s)+2(R(s-2,s-1)-1)+2r.$
Let N=rR(s2,s)+2R(s2,s1)+2rN=rR(s-2,s)+2R(s-2,s-1)+2r, $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$. For sake of contradiction, suppose there is no red $K_{s-1}$ nor blue $K_{s}$. If there is no blue edge in $G$, then we are done. So, set $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$. Without loss of generality, suppose $c(u,v)=B$. Then, $N(u)\cap N(v)\cong G^{\prime}=\textrm{KG}(N-2r,r)$. We will partition the vertices of $G^{\prime}$ into 4 sets. For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $w$ in $G^{\prime}$ such that $c(w,u)=X$ and $c(w,v)=Y$.
Observe that $N_{BR}$ must contain no red $K_{s-2}$ or we may find a red $K_{s-1}$ along with $v$, and $N_{BR}$ must contain no blue $K_{s-1}$ or we may find a blue $K_{s}$ along with $u$. Thus $\omega(G^{\prime}[N_{BR}])\leq R(s-2,s-1)-1$. Similarly, $N_{RR}\cup N_{RB}$ must contain no red $K_{s-2}$ or blue $K_{s}$, so $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq R(s-2,s)-1$. Furthermore, $N_{BB}$ must contain no red $K_{s-1}$ and no blue $K_{s-2}$, as otherwise we can find a blue $K_{s}$ along with the edge $uv$. Hence $\omega(G^{\prime}[N_{BB}])\leq R(s-2,s-1)-1$.
However, invoking Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(s-2,s)$), $F_{2}=N_{BR}$ ($k_{2}=R(s-2,s-1)$), and $F_{3}=N_{BB}$ ($k_{3}=R(s-2,s-1)$), since $k_{1}\cdot r+(k_{2}-1)+(k_{3}-1)=N-2r$, either $F_{1}$ contains a clique on R(s2,s)R(s-2,s) vertices, $F_{2}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, or $F_{3}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices. In any case, we are guaranteed either a red $K_{s-1}$ or a blue $K_{s}$, a contradiction. Therefore, $R^{\textrm{KG}}_{r}(s-1,s)\leq N$. ∎
Theorem 9 gives an improvement when $r=2$ and $s=3$.
Proof of Theorem 9.
For sake of contradiction, fix N>2(R(3,t2))+2tN>2(R(3,t-2))+2t, let $G=\textrm{KG}(N,2)$, $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$, and suppose there is no red $K_{3}$ or blue $K_{t}$. Note that if there are no blue edges, then there must be a red $K_{3}$, and we have arrived at our contradiction.
Hence, without loss of generality, we may assume $c(\{1,2\},\{3,4\})=B$. Next, we look at $N(\{1,2\})\cap N(\{3,4\})\cong G^{\prime}=\textrm{KG}(N-4,2)$, and we will partition the vertices of $G^{\prime}$ into $4$ sets. For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $v$ in $G^{\prime}$ such that $c(v,\{1,2\})=X$ and $c(v,\{3,4\})=Y$.
First, observe that for any two vertices $u,v\in N_{RB}$ with $uv\in E(G^{\prime})$, $c(u,v)=B$ (otherwise there is a red $K_{3}$ with $u,v,\{1,2\}$, a contradiction). Similarly, for each case of $u,v\in N_{RR}$, $u,v\in N_{BR}$, and $u\in N_{RR},v\in N_{RB}$, we have $c(u,v)=B$. These observations allow us to bound the clique numbers of $N_{RR}\cup N_{RB}$ and $N_{BR}$. In particular, $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq t-1$ since all edges are blue and we cannot have a blue $K_{t}$. Additionally, $\omega(G^{\prime}[N_{BR}])\leq t-2$ since all edges are blue, and if there were a blue $K_{t-1}$, together with vertex $\{1,2\}$, we would have a blue $K_{t}$.
Finally, $\omega(G^{\prime}[N_{BB}])<R(t-2,3)$ since otherwise, there would be either a copy of a red $K_{3}$ (a contradiction), or a blue copy of $K_{t-2}$, which together with the blue edge $(\{1,2\},\{3,4\})$ would be a blue $K_{t}$, another contradiction.
By Theorem 15, with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(t-2,3))$, $F_{2}=N_{BR}$ ($k_{2}=t$), and $F_{3}=N_{BB}$ ($k_{3}=t-1$), since $N-4\geq 2(R(t-2,3))+(t-1)+(t-1-1)$, either $F_{1}$ contains a $K_{R(t-2,3)}$, $F_{2}$ contains a $K_{t}$, or $F_{3}$ contains a $K_{t-1}$, a contradiction in all cases. This final contradiction proves the result. ∎
3.2 Lower bounds
The lower bound in Proposition 10 takes advantage of the chromatic number of Kneser graphs. Lovász  [5] proved that $\chi(\textrm{KG}(n,r))=n-2r+2$, which was one of the earliest applications of topological methods to combinatorics.
Proof of Proposition 10.
Let $n=R^{\textrm{KG}}_{r}(s,t)$ and set $G=\textrm{KG}(n,r)$. Suppose $\chi(G)=n-2r+2<R(s,t)$ and let $c:V(G)\to[n-2r+2]$ be a proper vertex-coloring of $G$. Consider the complete graph $G^{\prime}$ on n2r+2n-2r+2 vertices with $V(G^{\prime})=[n-2r+2]$ and let $C^{\prime}:E(G^{\prime})\to\{R,B\}$ be a red/blue edge-coloring of $G^{\prime}$ avoiding a red copy of $K_{s}$ and a blue copy of $K_{t}$.
Now, consider the following red/blue edge-coloring of $G$: For $uv\in E(G)$, color $uv$ with $C^{\prime}(c(u),c(v))$. We claim that this coloring has no red copy of $K_{s}$ or blue copy of $K_{t}$. Suppose there is a red copy of $K_{s}$ with vertices $\{v_{1},v_{2},\dots,v_{s}\}$. Observe that for all $i\neq j$, $c(v_{i})\neq c(v_{j})$ since $c$ is a proper vertex-coloring and $v_{i}v_{j}\in E(G)$. But now, the color of $v_{i}v_{j}$ is $C^{\prime}(c(v_{i}),c(v_{j}))=R$. This implies that the graph induced by the vertices $\{c(v_{1}),c(v_{2}),\dots,c(v_{s})\}$ is a red copy of $K_{s}$, a contradiction to our choice of $C^{\prime}$. Therefore, $n-2r+2\geq R(s,t)$. ∎
Proof of Theorem 11.
We wish to show there exists a red/blue coloring $c$ of $G=\textrm{KG}(3r+1,r)$ with no monochromatic triangle. To this end, we partition $V(G)$ as follows:
$X$ $=\{x\in V(G):3r+1\in x\},$
$Y$ $=\{x\in V(G)\setminus X:3r\in x\},$
$Z$ $=V(G)-Y-X\simeq V(\textrm{KG}(3r-1,r)).$
Observe that the induced subgraph $G[Z]$ of $G$ on $Z$ is triangle-free. Also, $X$ and $Y$ are independent sets with no triangles contained in $G[X\cup Y]$. Therefore, the only triangles will have either two vertices from $Z$ and one from $X\cup Y$ or one vertex from each of $X,Y,Z$. Setting $c(E(G[Z]))=B$ and $c(E(X,Y))=B$ with all other edges colored $R$ (see Figure 1), we obtain a red/blue coloring with no monochromatic triangles. ∎
XXYYZZ
Figure 1: The coloring in 11
3.3 Computationally-obtained bounds
To obtain the upper and lower bounds in Theorem 12 we wrote a program in Python, which for (small enough) specified parameters $n,r,s,t$, runs over all possible red/blue edge-colorings of $\textrm{KG}(n,r)$ and either provides as output a coloring that contains no red $K_{s}$ and no blue $K_{t}$ (which we call “good”), or declares that a good coloring does not exists.
We do this by reformulating the question as a satisfiability problem. For example, for $R^{\textrm{KG}}_{r}(3,3)$, we look for a red/blue edge-coloring $c$ of $G=\textrm{KG}(n,r)$ with no monochromatic triangle. Suppose $E(G)=\{e_{1},\dots,e_{|E(G)|}\}$. For $1\leq i\leq|E(G)|$, we may assign a boolean variable $v_{i}$ corresponding to $e_{i}$, where
$v_{i}=\begin{cases}1\text{ (True)}&\text{if }c(e_{i})=R,\\ 0\text{ (False)}&\text{if }c(e_{i})=B.\end{cases}$
Then, for a triangle $T=\{e_{1},e_{2},e_{3}\}$ of $G$, $T$ is monochromatic if and only if the following boolean clause is true:
$(v_{1}\land v_{2}\land v_{3})\lor(\neg v_{1}\land\neg v_{2}\land\neg v_{3}).$
Negating this, we get that $T$ is not monochromatic if and only if the following boolean clause, $B_{T}$, is true:
$B_{T}=(v_{1}\lor v_{2}\lor v_{3})\land(\neg v_{1}\lor\neg v_{2}\lor\neg v_{3}).$
Let $\{T_{1},\dots,T_{t}\}$ be the set of all triangles in $G$. Then, there is a good edge-coloring of $G$ if and only if there exists an assignment $(v_{1},\dots,v_{|E(G)|})\in\{0,1\}^{|E(G)|}$ such that the clause $B=\bigwedge_{i=1}^{t}B_{T_{i}}$ is satisfiable (True).
Our program builds the clause $B$ and then checks its satisfiability.
The proof of Theorem 12 follows because our program found a good coloring of $\textrm{KG}(8,2)$ (this coloring is given explicitly in Appendix A), but declared that the graphs $\textrm{KG}(9,2)$ and $\textrm{KG}(13,3)$ do not have good colorings.
Our code is given in Appendix B.
4 Pálvölgyi’s problem
We begin with a general discussion on monochromatic induced Kneser subgraphs, which will be useful in making progress on Problem 13.
Let $r>1$ be an integer and let $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ be Kneser graphs. Define the induced $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r,ind}(s,t)$ to be the minimum $n$ (if it exists) such that every red/blue edge-coloring of $\textrm{KG}(n,r)$ contains an induced red $\textrm{KG}(s,r)$ or an induced blue $\textrm{KG}(t,r)$.
Theorem 18.
For $r\geq 2$, $R^{\textrm{KG}}_{r,ind}(s,t)$ does not exists; that is, there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no red induced $\textrm{KG}(s,r)$ or blue induced $\textrm{KG}(t,r)$ .
In the proof of Theorem 18 we use the set-system theorem of Bollobás.
Theorem 19 (Bollobás [2]).
Let $(A_{1},B_{1}),\dots,(A_{m},B_{m})$ be pairs of sets with $|A_{i}|=a$ and $|B_{i}|=b$ for every $i$. Suppose that $A_{i}\cap B_{i}=\emptyset$ for $1\leq i\leq m$, and $A_{i}\cap B_{j}\neq\emptyset$ for $i\neq j$. Then $m\leq\binom{a+b}{a}$. Furthermore, if $m=\binom{a+b}{a}$ then there is some set $S$ of cardinality $a+b$ such that the sets $A_{i}$ are all the subsets of $S$ of size $a$, and $B_{i}=S\setminus A_{i}$ for all $i$.
Proof of Theorem 18.
We first show $R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist. Observe that $\textrm{KG}(2r,r)$ is a matching of size $\frac{1}{2}\binom{2r}{r}$.
Now, in any $\textrm{KG}(n,r)$ with $n\geq 2r$, we claim that every maximum induced matching $M$ has vertex set $\binom{S}{r}$, where $|S|=2r$. To this end, let $M$ be an induced matching with edge set $E=\{(X_{i},Y_{i})\colon 1\leq i\leq m\}$. Note that we must have $X_{i}\cap Y_{i}=\emptyset$ for all $i$, and $X_{i}\cap Y_{j}\neq\emptyset$ for all $i\neq j$ in order for $M$ to be an induced matching. Let $A_{i}=X_{i}$ for $1\leq i\leq m$, $A_{i}=Y_{i-m}$ for $m+1\leq i\leq 2m$, and let $B_{i}=Y_{i}$ for $1\leq i\leq m$, $B_{i}=X_{i-m}$ for $m+1\leq i\leq 2m$. Then the set pair system $(A_{i},B_{i})$, $1\leq i\leq 2m$ satisfies $A_{i}\cap B_{i}=\emptyset$ for all $i$ and $A_{i}\cap B_{j}\neq\emptyset$ for $i\neq j$.
Thus, we may invoke Theorem 19, with $a=b=r$, and
$2\cdot|E(M)|=2m\leq\binom{2r}{r}$.
In fact, since we know there exists an induced matching of size $\frac{1}{2}\binom{2r}{r}$, equality can be achieved, and there is a set $S$ such that $|S|=2r$ where $A_{i}\subseteq S$ and $B_{i}=S\setminus A_{i}$ for each $i$. Hence, one edge in a maximum induced matching $M$ uniquely determines the $2r$ elements of the ground set which can be used to form $M$.
Since a maximum induced matching $M$ can be uniquely determined by a single edge in $\textrm{KG}(n,r)$, this implies any two maximum induced matchings, i.e. matchings isomorphic to $\textrm{KG}(2r,r)$, are disjoint, and we can color each maximum induced matching bichromatically. Thus, we have a coloring of any $\textrm{KG}(n,r)$ with no monochromatic $\textrm{KG}(2r,r)$; i.e. $R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist.
Finally, observe that for any $s,t\geq 2r$, $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ each contain a subgraph isomorphic to $\textrm{KG}(2r,r)$. Hence, since there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no induced monochromatic $\textrm{KG}(2r,r)$ for any $n\geq 2r$, in the same edge-coloring there are no monochromatic induced $\textrm{KG}(s,r)$ nor monochromatic induced $\textrm{KG}(t,r)$. ∎
Proof of Theorem 14.
Set $n=R_{r}(3r,3r)-1$. We aim to show that there is a red/blue coloring $c$ of $G=\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle. We begin by partitioning $V(G)$ into two sets. Consider a red/blue edge coloring $c^{\prime}$ of $G^{\prime}=K^{r}_{n}$ containing no monochromatic $K^{r}_{3r}$. Observe that $E(G^{\prime})=V(G)$ and there is a clear bijection between the two sets. Let us partition $V(G)$ into two sets $P_{1},P_{2}$ as follows: For $e=\{v_{1},v_{2},\dots,v_{r}\}\in E(G^{\prime})$, if $c(e)=R$, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{1}$. Otherwise, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{2}$.
Observe that by the choice of $c^{\prime}$, neither $G[P_{1}]$ nor $G[P_{2}]$ contain an induced copy of $\textrm{KG}(3r,r)$ (since otherwise, $c^{\prime}$ would contain a monochromatic $K^{r}_{3r}$). Therefore, for any edge $e$ within $G[P_{1}]$ or $G[P_{2}]$, set $c(e)=R$ and otherwise, set $c(e)=B$. Since the blue subgraph is a bipartite graph, it is triangle-free. And so, $c$ is a red/blue edge coloring of $G$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle. ∎
5 Concluding remark
The multicolor Ramsey number $R(s_{1},s_{2},\dots,s_{t})$ is the minimum integer $n$ such that in any $t$-coloring of the edges of $K_{n}$, there exists a color-$i$ copy of $K_{s_{i}}$ for some $i\in[t]$. We can define the multicolor $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r}(s_{1},\dots,s_{t})$ analogously.
While we do not consider multicolor Kneser Ramsey numbers in depth in this paper, by similar arguments as in the proof of Theorem 17, we obtain an upper bound on these numbers.
Theorem 20.
Fix $s_{1},s_{2},\dots,s_{t}\geq 3$ and for $1\leq i\leq t$, set $m_{i}=R(s_{1},s_{2},\dots,s_{i}-1,\dots,s_{t})$. Suppose $m_{1}\geq m_{j}$ for all $j\geq 2$. Then,
$R^{\textrm{KG}}_{r}(s_{1},\dots,s_{t})\leq(m_{1}+1)r+\sum_{i=2}^{t}(m_{i}-1).$
Appendix A: A good coloring of $\textrm{KG}(8,2)$.
Here, we provide a coloring of $\textrm{KG}(8,2)$ with no monochromatic triangle.
The red edges are:
((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),\\ ((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),\\ ((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),\\ ((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),\\ ((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),\\ ((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),\\ ((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),\\ ((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),\\ ((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),\\ ((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),\\ ((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),\\ ((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),\\ ((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),\\ ((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),\\ ((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),\\ ((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),\\ ((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))
The remaining edges are blue.
We can describe part of this coloring as follows. First, we let $A=\binom{[3]}{2}$, $B=\binom{[8]-[3]}{2}$, and $C=V(G)-(A\cup B)$. Then all the edges from $A$ to $C$ as well as all the edges within $B$ are red, and all the edges from $A$ to $B$ are blue.
Appendix B: Our code
Below is the code we used to obtain the results in Theorem 12.
    import itertools as it
    from pysat.solvers import Glucose3
    Builds Kneser Graph Kn(n, r)
    Input: n, r
    Output: (V, E)
        V (list): Vertices (all r-subsets of [n]) as r-tuples
        E (list of 2-tuples): Edges as a pair of r-tuples
    def build_kneser_graph(n,r):
        V = list(it.combinations(range(1, n + 1), r))
        E = []
        for (a, b) in it.combinations(V, 2):
            if set(a).isdisjoint(set(b)):
                E.append((a, b))
        return (V, E)
    Try to find if Kn(n,r) can be 2-colored so that there is no red K_s or blue K_t
    For no red K_s, for every s-clique, need at least 1 blue edge.
    That is, if e_1, ..., e_k are the edges of an s-clique and e_i = 1 if red, e_i = -1 if blue,
    want to satisfy (-e_1 v -e_2 v ... v -e_k)
    For no blue K_t, for every t-clique, need at least 1 red edge.
    That is, if e_1, ..., e_k are the edges of a t-clique and e_i = 1 if red, e_i = -1 if blue,
    want to satisfy (-e_1 v -e_2 v ... v -e_k)
    (NOTE: 1 corresponds to True, -1 corresponds to False)
    If there is a good coloring, outputs (True, R, B), where R = red edges, B = blue edges
    If there is no good coloring, outputs (False, [], [])
    def build_sat_instances(n, r, s, t):
        solver = Glucose3()
        clauses = []
        V = []
        E = []
        #Create vertex and edge list
        (V, E) = build_kneser_graph(n, r)
        for s_tuple in it.combinations(V, s):
            #check that the s-tuple forms a clique in Kn(n,r)
            if all([set(s_tuple[i]).isdisjoint(set(s_tuple[j])) for (i,j) in it.combinations(range(len(s_tuple)), 2)]):
                #Add clause forbidding monochromatic K_s and K_t (if s = t)
                if s == t:
                    cliq = []
                    for edg in it.combinations(s_tuple, 2):
                        idx = -1
                        try:
                            idx = E.index(edg)
                        except:
                            idx = E.index((edg[1], edg[0]))
                        cliq.append(idx + 1)
                    solver.add_clause([-x for x in cliq])
                    solver.add_clause(cliq)
                #Add clause forbidding red K_s
                if s != t:
                    cliq = []
                    for edg in it.combinations(s_tuple, 2):
                        idx = -1
                        try:
                            idx = E.index(edg)
                        except:
                            idx = E.index((edg[1], edg[0]))
                        cliq.append(idx + 1)
                    solver.add_clause([-x for x in cliq])
        if s != t:
            for t_tuple in it.combinations(V, t):
                #check that the t-tuple forms a clique in Kn(n,r)
                if all([set(t_tuple[i]).isdisjoint(set(t_tuple[j])) for (i,j) in it.combinations(range(len(t_tuple)), 2)]):
                    #Add clause forbidding red K_t
                    cliq = []
                    for edg in it.combinations(t_tuple, 2):
                        idx = -1
                        try:
                            idx = E.index(edg)
                        except:
                            idx = E.index((edg[1], edg[0]))
                        cliq.append(idx + 1)
                    solver.add_clause(cliq)
        #Solve SAT problem and recover coloring
        result = solver.solve()
        R = []
        B = []
        if result:
            L = solver.get_model()
            R = [E[i - 1] for i in L if i > 0]
            B = [E[-i - 1] for i in L if i < 0]
        return (result, R, B)
    #output: True, R, B
    build_sat_instances(8, 2, 3, 3)
    #output: False, [], []
    build_sat_instances(9, 2, 3, 3)
    #output: False, [], []
    build_sat_instances(13, 3, 3, 3)